A ball is tossed straight up at 24m/s. Estimate its velocity at 1s interval, and when will it reach its original starting poin


1
Expert's answer
2022-04-04T09:03:21-0400

Given:

"v_0=24\:\rm m/s"

"g=9.8\:\rm m/s^2"

"t=1\:\rm s"


The velocity of a ball after 1s interval

"v=v_0-gt\\
=24-9.8*1=14.2 \:\rm m/s"